OpenStudy (anonymous):

what is the projection of (3,4) onto (3,1)

OpenStudy (zzr0ck3r):

\[projection\space a\rightarrow\space b=\frac{a\cdot b}{||b||}b\]

OpenStudy (zzr0ck3r):

what is \[u\cdot v=?\]

OpenStudy (anonymous):

thanks